OpenStudy (anonymous):

1 Using your function, explain to the Martians how to solve for f(3). 2. Describe to the Martians how to find the inverse of your function. 3. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2 My function is 4x+5=-3x+19

OpenStudy (mathmale):

If your function is 4x+5=-3x+19, the logical next step would be to solve it for x (not for f(3)). Add 3x to both sides and simplify the result. Then subtract 5 from both sides and simplify the result. What do you get?

OpenStudy (anonymous):

i did all that x=2 @mathmale

OpenStudy (mathmale):

Happy to hear that. Please share your work on the remaining problems, so I can see what you've already done and what you still need to do. Here you end up with 7x=14, so, yes, x=2. That leaves it something of a mystery what is meant by "solve for f(3)."

OpenStudy (mathmale):

What are your own thoughts on that?

OpenStudy (anonymous):

i need help solving for f(3)

OpenStudy (mathmale):

I understand that's what the question asks you to do. However, "your function" has neither f(x) nor y in i, so how are you going to evaluate f(3)?

OpenStudy (mathmale):

@johnlegend : OpenStudy tells me you're looking at someone else's math problem instead of at your own. When you're done, let me know if you want to finish your math problem.

OpenStudy (anonymous):

@mathmale i accidentally clicked it and was wondering why i was on it sorry @mathmale i didnt know i needed a y,because another person told me i didnt

OpenStudy (mathmale):

You are given the function / relationship / equation 4x+5=3x+19. This can be solved for x, as you have already done, but does not involve f(x) or y. However, we could CREATE OUR OWN FUNCTION, as follows: Rewrite 4x+5=3x+19 by moving the two terms on the right side to the left side AND THEN equating the result to f(x). Please do this. Type out what you now have.

OpenStudy (mathmale):

(Awaiting your response.)

OpenStudy (anonymous):

so would it be 4+5x=3+19x @mathmale

OpenStudy (mathmale):

I was asking you to start with that equation, but then to move the right two terms (the terms on the right side of your equation) to the left, and then labeling the result f(x). Please do that. 4 +5x = 3 + 19x -3-19x = -3 -19x ------- -------- combine. Set the resulting expression = to f(x).

OpenStudy (anonymous):

@mathmale are we solving for f(3) still?

OpenStudy (mathmale):

We will, but would you please finish this task first?

OpenStudy (anonymous):

oh im sorry 

OpenStudy (mathmale):



OpenStudy (anonymous):

no and im watching @mathmale

OpenStudy (mathmale):

Cool. Remember, we found that x=2 is a solution to your original equation. If we were to evaluate f(2), we'd get 7(2)-14=?? Try evaluating f(3).

OpenStudy (mathmale):

f(x)=7x-14. f(2)=7(2)-14 = ?

OpenStudy (mathmale):

Please go ahead. If your screen says "mathmale is typing a reply," ignore it and type your own reply.

OpenStudy (anonymous):

@mathmale 7 (3)-19 or is 7(3)-14 @mathmale

OpenStudy (mathmale):

Hi, John, what's up at your end?

OpenStudy (anonymous):

oh im sorry @mathmale you can continue

OpenStudy (mathmale):

John: f(x)=7x-14. If we evaluate this at x=2, we get f(2)=7(2)-14=14-14=0. If we evaluate this at x=3, we get what ??

OpenStudy (mathmale):

John, are we having technical problems? I'd like for this discussion to move faster, if at all possible.

OpenStudy (mathmale):

John: In the upper, left hand corner of your OpenStudy screen, you should see a symbol that looks like an envelope. This is for sending private messages. Send me a private message when you are ready to resume work on these problems that you've posted. In the meantime I'll be helping others.

OpenStudy (anonymous):

@mathmale openstudy seems to be moving slow on my computer so if you dont see me replying thats why but f(3)-14=21-14=7

OpenStudy (mathmale):

that's it. That's the answer you wanted. Let's review this.

OpenStudy (mathmale):

You were given an equation that had no f(x) or y in it. You were asked to evaluate f(3). This means that you had to create your own function from the given function. Rearranging the given relationship by moving the 2 terms on the right over to the left side, and setting the result equal to f(x), we got f(x)=7x-14.

OpenStudy (mathmale):

Note how f(2)=0. What this says, in effect, is, "Yes, 2 is a root of the original equation." Now you can evaluate f(x) for any x. Thus, f(3) = ?

OpenStudy (anonymous):

f(3)=7

OpenStudy (mathmale):

Right. And your function is f(x)=7x-14. How would you do Part 2 and find the inverse of this function? Have you done this type of problem before?

OpenStudy (anonymous):

Finding the inverse of an integer means performing an operation that will cancel out the number

OpenStudy (mathmale):

Let's paraphrase that: "Finding the inverse of a given FUNCTION results in obtaining a new FUNCTION that will cancel out the work of the original function."

OpenStudy (mathmale):

If f(x)=7x-14, replace f(x) with the label "y=" below:

OpenStudy (anonymous):

ok @mathmale what is y?

OpenStudy (mathmale):

I was just asking you to replace f(x) with y: f(x)=7x-14 => y=7x-14.

OpenStudy (mathmale):

Next, interchange x and y. Where you see x, write y; where you see y, write x. Result?

OpenStudy (anonymous):

ok x=2

OpenStudy (mathmale):

Take your y=7x-14. Replace that y with x and that x with y: x=7y-14. This is what I'd asked you to do.

OpenStudy (mathmale):

Now solve x=7y-14 for y. (This is a problem in algebra.)

OpenStudy (mathmale):

Step 1: Add 14 to both sides of the equation and simplify the result. What do you get? Type in your result.

OpenStudy (anonymous):

x=21y @mathmale

OpenStudy (mathmale):

Made a mistake; I'm sorry. Here's what I'd do: x = 7y - 14 +14= +14 --- ---------- Combine and simplify. Note: 7x + 14 does not equal 21 y. 7x and 14 are "unlike" terms and thus can't be combined.

OpenStudy (anonymous):

i didnt know you meant to all sides im sorry @mathmale so is it 14x=7y-0

OpenStudy (mathmale):

On the left side, x plus 14 equals what? Write this symbolically.

OpenStudy (anonymous):

14x @mathmale

OpenStudy (mathmale):

that looks like multiplication, not addition ("plus"). "x plus 14" could be written "14 added to x" and is represented symbolically by x+14. thus, on the left side of your equation you have just x+14; on the right side, you have 7y -14 + 14, or 7y + 0, or just 7y.

OpenStudy (mathmale):

In summary, we have x+14=7y, or 7y=x+14. How would you solve that for y?

OpenStudy (anonymous):

subtract 14 from both sides

OpenStudy (mathmale):

7y indicates multiplication. The operation opposite to multiplication is division. Therefore, to isolate y on the left side (to solve for y) requires that we DIVIDE each side by 7: \[7y=x+14\rightarrow \frac{ 7y }{ 7 }=\frac{ x }{ 7 }+\frac{ 14 }{ 7 }\]

OpenStudy (mathmale):

Please simplify this as best you can. Note again: We are trying to solve for y.

OpenStudy (mathmale):

Hint:\[\frac{ 7y }{ 7 }=y\]

OpenStudy (mathmale):

therefore, our equation becomes\[y=\frac{ x }{ 7 }+\frac{ 14 }{ 2 }\]If you were running out of time, you could leave this equation as is; it would represent the INVERSE FUNCTION of the given function. But go ahead and reduce 14/7. What's that?

OpenStudy (anonymous):

2

OpenStudy (mathmale):

Yes, so that leaves us with the equation of the inverse function, in simplest form:\[y=f ^{-1}(x)=\frac{ x }{ 7 }+2\]

OpenStudy (mathmale):

this completes the solution of Part 2. Questions? It's extremely important that you know how to do these operations yourself, so I'd strongly suggest that you review what we've done and then try to find the inverse of y=7x-14 yourself.

OpenStudy (mathmale):

@johnlegend : If you wish to continue, let me know what you want to do next.

OpenStudy (anonymous):

is x in the parentheses 2 @mathmale

OpenStudy (mathmale):

Not sure I understand your question. Your function, f (x), is 7x-14. Your inverse function, the inverse of f(x)=7x-14, is \[f ^{-1}(x)=\frac{ x }{ 7 }+2\] What are you asking?

OpenStudy (anonymous):

do we have to solve the inverse function @mathmale

OpenStudy (mathmale):

for Part 2, our job was to find the inverse function of f(x)=7x-14. We did. See the formula for the inverse function, above and below:\[f ^{-1}(x)=\frac{ x }{ 7 }+2\] So, there's nothing to "solve." What next?

OpenStudy (anonymous):

part 3 @mathmale